At the start of 2020, the federal debt is 21 trillion. During 2020, the government spends 3.5 trillion, and takes in 2.5 trillion in taxes.
        
What will the federal debt be at the start of 2021?
Answer #1

Solution-

Federal debt during 2020=

Total spending- Total earnings

(Total revenue = tax = 2.5 trillion)

=3.5-2.5= 1 trillion.

So, Federal debit at the start of 2021=

Federal debit at the start of 2020 + Federal debit during 2020

= 21+1 =22 trillion.

​​​​​​